OpenStudy (anonymous):

5a + 5 = 35 Evaluate the equation with a given solution set.

OpenStudy (anonymous):

first subtract 5 from both sides.

OpenStudy (unklerhaukus):

...then divide both sides by 5

OpenStudy (anonymous):

ok I eill do that

OpenStudy (anonymous):

Got 30 and 0

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

35 - 5 = 30 5 - 5 = 0 30 ÷ 5 = 6 0 ÷ 5 = 0 a= 6

razor99 (razor99):

5a+5=35 5a=35-5 5a=30 a=30/5 a=6

OpenStudy (anonymous):

THanks! :D

razor99 (razor99):

no problem :D